OpenStudy (anonymous):

Describe the steps you would take to write an equation in point-slope form of the line that passes through the points (3,-2) and (4,5).

OpenStudy (mertsj):

slope = 5-(-2)/4-3 = 7/1=7 so choose either point and write y-5 = 7(x-4)

OpenStudy (anonymous):

Describe the steps*